# A modern perspective on rational homotopy theory

### v1

@@ 2.5. The $1$st Eilenberg subcomplex of a pointed simplicial set
This universal property amounts to the promised adjunction, which is stated in Part (2) of the following lemma.
Lemma 2.23.
@@
    The $1$st Eilenberg subcomplex functor $E_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is right adjoint to $F$.
Lemma 2.23 implies that small limits and colimits in ${\mathcal{SS}\mathrm{ets}}_{1}$ are constructed as in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. Given a diagram
$L\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$
in $1$-reduced simplicial sets, we form its (co)limit by post-composing $L$ with the forgetful functor
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
and forming the (co)limit of the diagram
$FL\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
in pointed simplicial sets. This (co)limit is a $1$-reduced simplicial set, and all the structure maps are in ${\mathcal{SS}\mathrm{ets}}_{1}$.
In particular, the initial object of ${\mathcal{SS}\mathrm{ets}}_{1}$ is the point $\Delta[0]$, not the empty simplicial set $\emptyset$. Thus, the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$ does not preserve the initial object, so it cannot admit a right adjoint.
@@ 3. The left transfer to $1$-reduced simplicial sets
We transfer the model structure on the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets along the adjunction
@@
to a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. This left transfer is an application of [BHK+15, Theorem 2.23].
@@ 3.1. Model categories
We collect some background on model categories from [DS95, §3] and [Hir03, Chapters 7, 11, and 13]. We first recall a concept in the homotopy theory of simplicial sets from [GJ99, Chapter I].
@@
A Kan fibration is a simplicial map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the set of all horn inclusions
$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
In other words, a simplicial map $p\colon X\rightarrow Y$ is a Kan fibration if every lifting problem of the form
${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$${Y}$$\scriptstyle{j_{n,k}}$$\scriptstyle{p}$
@@
Example 3.2 ([GJ99, Chapter I, Lemma 3.3]).
@@
The homotopy theory of simplicial sets can be studied systematically in the framework of model categories.
@@
A model category comprises a category $\mathcal{M}$ that has all finite limits and colimits, as well as three distinguished classes of morphisms in $\mathcal{M}$: a class of weak equivalences, a class of cofibrations, and a class of fibrations. A weak equivalence that is also a cofibration is called an acyclic cofibration. A weak equivalence that is also a fibration is called an acyclic fibration. This data satisfies the axioms in [DS95, Definition 3.3].
@@
In a model category, the weak equivalences and the fibrations together determine the cofibrations. Dually, the weak equivalences and the cofibrations together determine the fibrations.
Remark 3.5.
Let $\mathcal{M}$ be a model category with initial object $\emptyset$ and terminal object $\ast$. An object $X$ of $\mathcal{M}$ is
  1. (1)
    cofibrant if the unique morphism $\emptyset\rightarrow X$ is a cofibration; and
  2. (2)
    fibrant if the unique morphism $X\rightarrow\ast$ is a fibration.
Then, for an object $Y$ of $\mathcal{M}$,
  1. (1)
    a cofibrant replacement of $Y$ is a weak equivalence $A\rightarrow Y$ with cofibrant source $A$; and
  2. (2)
    a fibrant replacement of $Y$ is a weak equivalence $Y\rightarrow B$ with fibrant target $B$.
Example 3.6 ([Hir03, Theorems 7.10.12 and 7.10.13]).
@@
    a weak equivalence if and only if $f$ is a weak homotopy equivalence;
@@
The model categories ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ have the additional property of being cofibrantly generated. In practice, the property of being cofibrantly generated facilitates the detection of (acyclic) fibrations in a model category.
Definition 3.7 ([Hir03, Definition 11.1.2]).
@@
    both $I$ and $J$ permit the small object argument [Hir03, Definition 10.5.15];
@@
    $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}.$
    Attaching a disjoint basepoint yields a generating set of cofibrations $I_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
@@
    $J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
    Attaching a disjoint basepoint yields a generating set of acyclic cofibrations $J_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
@@
Definition 3.9 ([Dug01, Definition 2.1]).
A model category $\mathcal{M}$ is combinatorial if $\mathcal{M}$ is cofibrantly generated and its underlying category is locally presentable [AR94, Definition 1.17].
Example 3.10 ([AR94, Example 1.12]).
@@
Lastly, ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ satisfy the following model-categorical property that permits left Bousfield localization.
Definition 3.11 ([Hir03, Definition 13.1.1(1)]).
@@
The following result guarantees the left properness of almost every model category that appears in this paper.
Proposition 3.12 ([Hir03, Corollary 13.1.3(1)]).
@@
3.2. The left transfer to $1$-reduced simplicial sets
Firstly, to appreciate the necessity of using pointed simplicial sets and their $1$st Eilenberg subcomplex, we explain why we cannot use the $1$-reduction functor $R_{1}$ to right-induce, in the sense of [Hir03, Theorem 11.3.2], a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ along the adjunction
@@ 3.2. The left transfer to $1$-reduced simplicial sets
from Lemma 2.19. We write $J$ for the set of all horn inclusions in ${\mathcal{SS}\mathrm{ets}}$. If such a right transfer were possible, then, in the resulting right-induced model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$,
@@
Thus, every map in $R_{1}(J)$ would be a weak homotopy equivalence. We verify that this necessary condition fails.
@@
    If $n=1$, then ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point.
@@
    If $n=2$, then ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks out the unique 00-simplex of the simplicial $2$-sphere.
@@
    If $n\geq 3$, then $|{R}_{1}\left(j_{n,k}\right)|$ is the inclusion of a strong deformation retract in topological spaces.
Therefore, ${R}_{1}\left(j_{n,k}\right)$ is a weak homotopy equivalence if and only if $n\neq 2$.
@@
Parts (1) and (2) are straightforward. For (3), if $n\geq 3$, then $\operatorname{sk}_{1}\left(\Lambda[n,k]\right)=\operatorname{sk}_{1}\left(\Delta[n]\right)$, so the inclusion of a strong deformation retract $|j_{n,k}|$ descends to the inclusion of a strong deformation retract $|{R}_{1}\left(j_{n,k}\right)|$. ∎
@@
We left-induce a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ by transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ along the adjunction
@@
from Lemma 2.23(2). To do so, we apply the following general result on left-induced model structures.
Theorem 3.15 ([BHK+15, Theorem 2.23]).
Let $F\colon\mathcal{C}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{C}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $Q$. Suppose that the following condition holds: if a morphism $g$ in $\mathcal{C}$ has the right lifting property with respect to the class $F^{-1}(Q)$, then $F(g)$ is a weak equivalence in $\mathcal{M}$. Then, there is a combinatorial model structure on $\mathcal{C}$ in which a morphism $f$ is
@@
As ${\mathcal{SS}\mathrm{ets}}_{1}$ is a locally presentable category [AR94, Example 1.12] and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is a combinatorial model category by Example 3.10, it remains to verify the condition in Theorem 3.15.
@@
Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is a weak homotopy equivalence.
@@
We show that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$. For $n\geq 0$, let $i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$ denote the boundary inclusion in ${\mathcal{SS}\mathrm{ets}}$. By Lemma 3.14, ${R}_{1}(i_{n})$ is a monomorphism of $1$-reduced simplicial sets. Therefore, the hypothesis informs us that, for every $n\geq 0$, $f$ has the right lifting property with respect to ${R}_{1}(i_{n})$. Then, the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19 yields that, for every $n\geq 0$, $f$ has the right lifting property with respect to the boundary inclusion $i_{n}$. We conclude that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$. ∎
We begin our recovery of the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2] and the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6].
@@
    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
@@
${{\Lambda[n,k]}}$${X}$${{\Delta[n].}}$$\scriptstyle{j_{n,k}}$
Using the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we may equivalently solve a lifting problem
${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$
In light of Lemma 3.13, we split our solution in two separate cases. If $n\neq 2$, then, by Lemmas 3.13 and 3.14, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, so our lifting problem has a solution since $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$. If $n=2$, then we solve a lifting problem
@@
where $x_{0}$ denotes the unique 00-simplex of $X$, by setting $\psi$ to be the constant map at $x_{0}$. ∎
The left transfer in Theorem 3.17 renders the adjunction
from Lemma 2.23(2) a Quillen adjunction; in particular, we note that $E_{1}$ preserves fibrant objects.
Corollary 3.18.
If $X$ is a pointed Kan complex, then $E_{1}(X)$ is a Kan complex.
In particular, we occasionally appeal to a special instance of Corollary 3.18.
@@
Proof.
By Example 3.2, $\operatorname{Sing}(Y)$ is a pointed Kan complex, so the claim follows from Corollary 3.18. ∎

### v2

@@ 2.5. The 11st Eilenberg subcomplex of a pointed simplicial set
This universal property amounts to the promised adjunction, which is in part (2) of the following lemma.
Proposition 2.19.
@@
    The 11st Eilenberg subcomplex functor $E_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is right adjoint to $F$.
Proposition 2.19 implies that small limits and colimits in ${\mathcal{SS}\mathrm{ets}}_{1}$ are constructed as in $\mathcal{SS}\mathrm{ets}$$_{\ast}$. Given a diagram $L\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ in $1$-reduced simplicial sets, we form its (co)limit by post-composing $L$ with the forgetful functor $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ and forming the (co)limit of the diagram $FL\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ in pointed simplicial sets. In particular, coproducts in ${\mathcal{SS}\mathrm{ets}}_{1}$ are given by wedge sums instead of disjoint unions, for the latter do not yield a $1$-reduced simplicial set. More fundamentally, the initial object of ${\mathcal{SS}\mathrm{ets}}_{1}$ is the point $\Delta[0]$; the empty set $\emptyset$ is not $1$-reduced. Thus, the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$ does not preserve the initial object or coproducts, so it cannot admit a right adjoint.
@@ 3. The left transfer to $1$-reduced simplicial sets
We recover the first model structure on $1$-reduced simplicial sets from Quillen’s family, which we then localize to recover the rest of the family. Namely, we left-transfer the model structure on the category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ of pointed simplicial sets along the adjunction
@@
to a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. (Our notation for adjunctions is amenable to a helpful mnemonic that distinguishes a left transfer from a right transfer; we propagate the existence of a model structure from right to left, so it is a left transfer.) This left transfer is an application of [4, Theorem 2.23], which we recall after going over the background on model category theory that we rely on in this paper.
@@ 3.1. Model categories
We collect some background on model categories from [18, §3] and [26, Chapters 7, 11, and 13]. We first recall a concept in the homotopy theory of simplicial sets from [23, Chapter I].
@@
A Kan fibration is a simplicial map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the set of all horn inclusions $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}$. In other words, a simplicial map $p\colon X\rightarrow Y$ is a Kan fibration if every lifting problem of the form
[…]
@@
Example 3.2 ([23, Chapter I, Lemma 3.3]).
@@
Example 3.2 follows from the geometric realization-singular complex adjunction combined with the fact that, at the level of topological spaces, geometric realizations of horns can always be filled because they are retracts of geometric realizations of standard simplices. However, the directionality of simplices in a simplicial set precludes the existence of the analogous retraction at the level of simplicial sets, so the concepts of a Kan fibration and a Kan complex are not moot.
The homotopy theory of simplicial sets can be studied systematically in the framework of model categories, which is the framework that we adhere to throughout this paper.
@@
A model category comprises a category $\mathcal{M}$ that has all finite limits and colimits, as well as three distinguished classes of morphisms in $\mathcal{M}$: a class of weak equivalences, a class of cofibrations, and a class of fibrations. A weak equivalence that is also a cofibration is called an acyclic cofibration. A weak equivalence that is also a fibration is called an acyclic fibration. This data satisfies the axioms in [18, Definition 3.3].
@@
In a model category, the weak equivalences and the cofibrations together determine the fibrations.
To recover Quillen’s family of model structures on $1$-reduced simplicial sets, we rely on the standard model structure on (pointed) simplicial sets.
Example 3.5 ([26, Theorems 7.10.12 and 7.10.13]).
@@
    a weak equivalence if and only if $|f|$ is a weak homotopy equivalence;
@@
The following definition is intended to introduce some terminology for the rest of this paper.
Definition 3.6.
Let $\mathcal{M}$ be a model category with initial object $\emptyset$ and terminal object $\ast$. An object $X$ of $\mathcal{M}$ is cofibrant if the morphism $\emptyset\rightarrow X$ is a cofibration, and $X$ is fibrant if the morphism $X\rightarrow\ast$ is a fibration. Then, for an object $Y$ of $\mathcal{M}$, a cofibrant replacement of $Y$ is a weak equivalence $A\rightarrow Y$ with cofibrant source $A$, and a fibrant replacement of $Y$ is a weak equivalence $Y\rightarrow B$ with fibrant target $B$.
The model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ have the property of being cofibrantly generated. In practice, this property facilitates the detection of (acyclic) fibrations and fibrant objects in a model category.
Definition 3.7 ([26, Definition 11.1.2]).
@@
    both $I$ and $J$ permit the small object argument [26, Definition 10.5.15];
@@
    $I=\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\}.$
@@
    $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}.$
Attaching a disjoint basepoint to the maps in both sets yields the generating (acyclic) cofibrations of $\mathcal{SS}\mathrm{ets}$$_{\ast}$.
@@
Definition 3.9 ([17, Definition 2.1]).
A model category $\mathcal{M}$ is combinatorial if $\mathcal{M}$ is cofibrantly generated and its underlying category is locally presentable [1, Definition 1.17].
In a combinatorial model category, not only do we have a set of building blocks for its (acyclic) cofibrations by virtue of cofibrant generation, but we also have a set of building blocks for its objects by virtue of local presentability. Though the category of topological spaces is not locally presentable [1, Example 1.18(5)], the category of simplicial sets does not run into this problem as every simplicial set is built out of standard simplices.
Example 3.10 ([1, Example 1.12]).
@@
Also, $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ have the following model-categorical property that permits left Bousfield localization.
Definition 3.11 ([26, Definition 13.1.1(1)]).
@@
The following result guarantees the left properness of almost every model category that appears in this paper, including every model structure in Quillen’s family of model structures on $1$-reduced simplicial sets.
Proposition 3.12 ([26, Corollary 13.1.3(1)]).
@@
In practice, it is common and convenient for the cofibrations in a model category to be the monomorphisms, as is the theme in this paper. Then, every object is cofibrant, so we need not worry about cofibrant replacement, and we can use Proposition 3.12 to extract left properness for free.
3.2. The left transfer
Before we do the promised left transfer, we explain why we cannot use the $1$-reduction functor $R_{1}$ to right-induce, in the sense of [26, Theorem 11.3.2], a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ along the adjunction
@@ 3.2. The left transfer
from Proposition 2.15. We write $J$ for the set of all horn inclusions in $\mathcal{SS}\mathrm{ets}$}. If such a right transfer were possible, then, in the resulting right-induced model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$,
@@
In particular, the set $|R_{1}(J)|$ would comprise weak homotopy equivalences, which is false by part (2) of the following lemma; conceptually, we find an instance of the $1$-reduction quotient being homotopically ill-behaved.
@@
    If $n=1$, then ${R}_{1}(j_{1,k})$ is the identity map of the point.
@@
    If $n=2$, then ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks out the unique 0-simplex of the simplicial $2$-sphere.
@@
    If $n\geq 3$, then $|{R}_{1}(j_{n,k})|$ is the inclusion of a strong deformation retract in topological spaces.
Thus, $|{R}_{1}(j_{n,k})|$ is a weak homotopy equivalence if and only if $n\neq 2$.
@@
Parts (1) and (2) are straightforward. For (3), if $n\geq 3$, then $\operatorname{sk}_{1}(\Lambda[n,k])=\operatorname{sk}_{1}(\Delta[n])$, so the inclusion of a strong deformation retract $|j_{n,k}|$ descends to the inclusion of a strong deformation retract $|{R}_{1}(j_{n,k})|$. ∎
@@
Instead, we left-induce a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ by transferring that on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ along the adjunction
@@
from Proposition 2.19(2). To do so, we apply the following general result on left-induced model structures.
Theorem 3.15 ([4, Theorem 2.23]).
Let $F\colon\mathcal{C}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{C}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $Q$. Suppose that, if a map $g$ in $\mathcal{C}$ has the right lifting property with respect to the class $F^{-1}(Q)$, then $F(g)$ is a weak equivalence in $\mathcal{M}$. Then, there is a combinatorial model structure on $\mathcal{C}$, called the left-induced model structure, in which a map $f$ is
@@
As ${\mathcal{SS}\mathrm{ets}}_{1}$ is a locally presentable category [1, Example 1.12] and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ is a combinatorial model category by Example 3.10, it remains to verify that our adjunction satisfies the homotopical condition in Theorem 3.15.
@@
Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $|f|$ is a weak homotopy equivalence.
@@
We show that $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}. For $n\geq 0$, let $i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$ denote the boundary inclusion in $\mathcal{SS}\mathrm{ets}$}. By Lemma 3.14, ${R}_{1}(i_{n})$ is a monomorphism of $1$-reduced simplicial sets. Therefore, the hypothesis informs us that, for every $n\geq 0$, $f$ has the right lifting property with respect to ${R}_{1}(i_{n})$. Then, the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15 yields that, for every $n\geq 0$, $f$ has the right lifting property with respect to the boundary inclusion $i_{n}$. We conclude that $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}. ∎
We begin recovering the family of model structures on $1$-reduced simplicial sets in [33, Part II, Theorem 2.2].
@@
    a weak equivalence if and only if $|f|$ is a weak homotopy equivalence; and
@@
Using the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, we equivalently solve a lifting problem
[…]
In light of Lemma 3.13, we split our solution in two separate cases. If $n\neq 2$, then, by Lemmas 3.13 and 3.14, ${R}_{1}(j_{n,k})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, so our lifting problem has a solution since $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$. If $n=2$, then we solve a lifting problem
[…]
where $x_{0}$ denotes the unique 0-simplex of $X$, by setting $\psi$ to be the constant map at $x_{0}$. ∎
The left transfer renders $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$ a Quillen adjunction [26, Definition 8.5.2], so $E_{1}$ preserves fibrant objects; an alternative proof of this preservation uses the pullback definition of the 11st Eilenberg subcomplex together with the fact that, if $X$ is a Kan complex, then $X\rightarrow\operatorname{cosk}_{1}(X)$ is a Kan fibration. We record the following consequence for the next section.